The real interest rate is 7%. Ken was offered two possible compensation arrangements by his boss: option A where y1=35,000 and y2=36,000; option B where y1=36,000 and y2=34,900. Ken prefers option A, option B, indifferent [choose one]. Internal Revenue Service offers Ken two possible tax payment arrangements: option C where t1=4000, t2=4000; option D where t1=2000 and t2=6100. Ken prefers option C, option D, indifferent [choose one]

Just show the second blank process. (option C or option D)
Answer #1

1)

Option A

Compensation in period 1=$35000

Compensation in period 2=$36000

PW of compensations=35000+36000/(1+7%)=$68644.86

Option B

Compensation in period 1=$36000

Compensation in period 2=$34900

PW of compensations=36000+34900/(1+7%)=$68616.82

Option A gives higher present worth. Ken would prefer option A

2)

Tax arrangement C

Tax in period 1=4000

Tax in period 2=4000

PW of tax=4000+4000/(1+7%)=$7738.32

Tax arrangement D

Tax in period 1=2000

Tax in period 2=6000

PW of tax=2000+6000/(1+7%)=$7607.48

PW of tax is lower in case of arrangement D. So, Ken would prefer option D